# USimAgent: Large Language Models for Simulating Search Users

### v1

@@ Abstract.
Due to the advantages in the cost-efficiency and reproducibility, user simulation has become a promising solution to the user-centric evaluation of information retrieval systems. Nonetheless, accurately simulating user search behaviors has long been a challenge, because users’ actions in search are highly complex and driven by intricate cognitive processes such as learning, reasoning, and planning. Recently, Large Language Models (LLMs) have demonstrated remarked potential in simulating human-level intelligence and have been used in building autonomous agents for various tasks. However, the potential of using LLMs in simulating search behaviors has not yet been fully explored. In this paper, we introduce a LLM-based user search behavior simulator, USimAgent. The proposed simulator can simulate users’ querying, clicking, and stopping behaviors during search, and thus, is capable of generating complete search sessions for specific search tasks. Empirical investigation on a real user behavior dataset shows that the proposed simulator outperforms existing methods in query generation and is comparable to traditional methods in predicting user clicks and stopping behaviors. These results not only validate the effectiveness of using LLMs for user simulation but also shed light on the development of a more robust and generic user simulators.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Recently, user simulation has become a promising solution to the user-centric evaluation of information retrieval systems (Balog and Zhai, 2023). By generating a vast amount of data through simulators without the need for actual user involvement, user simulation offers not only cost-effectiveness but also significant advantages in terms of flexibility and efficiency in data collection. By adjusting simulation rules, data structures can be conveniently added or modified to swiftly adapt to changing research requirements. Additionally, unlike human participants, simulators are not prone to issues like fatigue from prolonged search activities, ensuring the consistency and repeatability of experimental results.
Given the complexity of user search behavior, simulating the whole search session has long been a challenging and enduring task. Traditional simulation methods typically decompose a user’s interactive search behavior into a series of independent steps, including submitting queries, browsing Search Engine Results Pages (SERPs), clicking results, reading and evaluating documents, and deciding when to stop (Balog and Zhai, 2023). Therefore, they require a dedicated simulation strategy to be designed for each step, such as generating search queries by extracting terms from language models associated with specific documents or topics(Azzopardi et al., 2007; Azzopardi, 2009; Jordan et al., 2006; Baskaya et al., 2012; Carterette et al., 2015), estimating the probability of a user clicking on search results based on historical data using click models(Craswell et al., 2008; Guo et al., 2009; Chapelle and Zhang, 2009; Dupret and Piwowarski, 2008; Borisov et al., 2016), and deciding when a user stops searching based on a set of predefined simplistic assumptions through heuristic rules(Cooper, 1973; Kraft and Lee, 1979; Nickles et al., 1995; Maxwell et al., 2015; Maxwell, 2019). However, these approaches fail to fully consider the dynamic and interdependent nature of user behavior, In particular, they often fail to model how some cognitive factors, such as the information needs and background knowledge, and cognitive processes, including the learning and reasoning, would affect and drive user’s action. These limitations make it difficult for the existing user simulation methods to accurately and comprehensively reproduce the complex interactions between users and search engines.
@@ 3. METHODOLOGY
(1) $I\leftarrow\begin{cases}I&\text{if }s_{t}=\text{continue}\\ (I,s)&\text{if }s_{t}=\text{stop}\end{cases}$
When $s_{t}$ is to continue searching, it moves to the query reformulation step; otherwise, the session ends, and the action sequence $I$ is output.
@@ 4.1.1. Datasets
We evaluated our USimAgent on public user behavior dataset, which was collated by Liu et al. through a laboratory user study(Liu et al., 2019). Participants completed nine intricate search tasks, with the dataset incorporating their search interactions, which include query reformulations, clicks, and stopping behaviors.
@@ 4.1.3. Implementation Details
Our system employs GPT-4333https://chat.openai.com/, a widely recognized and extensively utilized LLM, with the configuration settings of “temperature = 0, n = 1”.

### v2

@@ Abstract.
Due to the advantages in the cost-efficiency and reproducibility, user simulation has become a promising solution to the user-centric evaluation of information retrieval systems. Nonetheless, accurately simulating user search behaviors has long been a challenge, because users’ actions in search are highly complex and driven by intricate cognitive processes such as learning, reasoning, and planning. Recently, Large Language Models (LLMs) have demonstrated remarked potential in simulating human-level intelligence and have been used in building autonomous agents for various tasks. However, the potential of using LLMs in simulating search behaviors has not yet been fully explored. In this paper, we introduce a LLM-based user search behavior simulator, USimAgent. The proposed simulator can simulate users’ querying, clicking, and stopping behaviors during search, and thus, is capable of generating complete search sessions for specific search tasks. Empirical investigation on a real user behavior dataset shows that the proposed simulator outperforms existing methods in query generation and is comparable to traditional methods in predicting user clicks and stopping behaviors. These results not only validate the effectiveness of using LLMs for user simulation but also shed light on the development of a more robust and generic user simulators. The code and data are accessible at https://github.com/Meow-E/USimAgent.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657963isbn: 979-8-4007-0431-4/24/07ccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Recently, user simulation has become a promising solution to the user-centric evaluation of information retrieval systems (Balog and Zhai, 2023). By generating a vast amount of data through simulators without the need for actual user involvement, user simulation offers not only cost-effectiveness but also significant advantages in terms of flexibility and efficiency in data collection. Additionally, simulators are not prone to issues like fatigue from prolonged search activities, ensuring the consistency and repeatability of experimental results.
Given the complexity of user search behavior, simulating the whole search session has long been a challenging and enduring task. Traditional simulation methods typically decompose a user’s interactive search behavior into a series of independent steps, including submitting queries, browsing Search Engine Results Pages (SERPs), clicking results, reading and evaluating documents, and deciding when to stop (Balog and Zhai, 2023). Therefore, they require a dedicated simulation strategy to be designed for each step, such as generating search queries by extracting terms from language models associated with specific documents or topics(Azzopardi et al., 2007; Azzopardi, 2009; Jordan et al., 2006; Baskaya et al., 2012; Carterette et al., 2015), estimating the probability of a user clicking on search results based on historical data using click models(Craswell et al., 2008; Guo et al., 2009; Chapelle and Zhang, 2009; Dupret and Piwowarski, 2008; Borisov et al., 2016), and deciding when a user stops searching based on a set of predefined simplistic assumptions through heuristic rules(Cooper, 1973; Kraft and Lee, 1979; Nickles et al., 1995; Maxwell et al., 2015; Maxwell, 2019). However, these approaches fail to fully consider the dynamic and interdependent nature of user behavior. In particular, they often fail to model how some cognitive factors, such as the information needs and background knowledge, and cognitive processes, including the learning and reasoning, would affect and drive user’s action. These limitations make it difficult for the existing user simulation methods to accurately and comprehensively reproduce the complex interactions between users and search engines.
@@ 3. METHODOLOGY
(1) $\ I\leftarrow\begin{cases}I&\text{if }s_{t}=\text{continue}\\ (I,s)&\text{if }s_{t}=\text{stop}\end{cases}\vspace{-5pt}$
When $s_{t}=\text{continue}$, it moves to the query reformulation step; otherwise, the session ends, and the action sequence $I$ is output.
@@ 4.1.1. Datasets
We evaluated our USimAgent on public user behavior dataset [^1] http://www.thuir.cn/KDD19-UserStudyDataset/, which was collated by Liu et al. through a laboratory user study(Liu et al., 2019). Participants completed nine intricate search tasks, with the dataset incorporating their search interactions, which include query reformulations, clicks, and stopping behaviors.
@@ 4.1.3. Implementation Details
Our system employs GPT-$4$https://chat.openai.com/, a widely recognized and extensively utilized LLM, with the configuration settings of “temperature = 0, n = 1”.
@@ 5. Conclusion and future work
Acknowledgements.
This research was supported by the Natural Science Foundation of China (61902209, 62377044, U2001212), and Beijing Outstanding Young Scientist Program (NO. BJJWZYJH012019100020098), Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the ”Double-First Class” Initiative, Renmin University of China.